Selected year-end financial statements of Cabot Corporation follow. (All sales were on credit selected balance sheet amounts at December 31, 2016, were Inventory, $53,900; total assets, $239,400; common stock, $87,000; and retalned earnings, $35104.) CABOT CORPORATION Income Statement For Year Ended December 31, 2017 448, 600 298,150 150,450 99,000 4, 400 47,050 18,954 28 , 096 ales $ Cost of goods sold Gross profit Income before taxes Income taxes Net income CABOT CORPORATION Balance Sheet December 31, 2017 Assets Cash hort-term investments Accounts receivable, net Notes receivable (trade) Merchandise inventory Liabilities and Equity $ 14, 000 Accounts payable $ 19,500 3,400 2, 900 B,200 crued wages payable 2B, B00 Income taxes payable 5,000 36,150 Long-term note payable, secured by 71,400 mortgage on plant assets Prepaid expenses Plant assets, net 2, 950 Coamon stoch 152,300 Retained earnings B1,000 63, 200 $ 247,400 Total assets $ 247,400 Total liabilities and equity These are short-term notes recelvable arising from customer (trade) sales. Required: Compute the following: (1) current ratio, (2) acld-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory. 6) debt-to-equity ratlo, (7) times Interest earned, (8) profit margin ratio. (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equity. (Do not round Intermediate calculations.) Complete this question by entering your answers in the tabs below. Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11mortgage on plant assets Prepaid expenses Plant assets, net 2, 950 Coamon stoch 152, 300 Retained earnings B1,000 63, 200 $ 247,400 Total assets $ 247.400 Total liabilities and equity These are short-term notes recelvable arising from customer (trade) sales. Required: Compute the following: (1) current ratio, (2) acld-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory. (6) debt-to-equity ratlo, (7) times Interest earned, (8) profit margln ratio. (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equity. (Do not round intermediate calculations.) Complete this question by entering your answers in the tabs below. Req 1 and 2Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the days sales uncollected. Sales Uncoll Choose Numerator: x DaysDays Sales Uncollected Accounts Receivable net (includingrntNet sales x | 365 | | Days sales uncollected customers) 2017 O days < Req 1 and 2 Req 4These are short-term notes recelvable arising from customer (trade) sales. Required: Compute the following: (1) current ratio, (2) acld-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory. (6) debt-to-equity ratlo, (7) times Interest earned, (8) profit margln ratio. (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equity. (Do not round intermediate calculations.) Complete this question by entering your answers in the tabs below. Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the inventory turnover. Choose Numerator: Cost of goods sold Average inventory - Inventory turnover 0 times 2017 K Req 3 Req 5Notes receivable (trade) 5,000 36,150 Long-tern note payable, secured by 71,400 mortgage on plant assets Prepaid expenses Plant assets, net 2, 950 Coamon stoch 152,300 Retained earnings B1,000 63, 200 $ 247,400 $ 247.400 Total liabilities and equity These are short-term notes recelvable arising from customer (trade) sales. Required: Compute the following: (1) current ratio, (2) acld-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory. (6) debt-to-equity ratlo, (7) times Interest earned, (8) profit margln ratio. (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equity. (Do not round intermediate calculations.) Complete this question by entering your answers in the tabs below. Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the days sales in inventory. Choose Choose Numerator: x DaysDays Sales in Inventory inventoryCost of goods sold 305 | | Days sales in inventory x = 2017 0 days K Req 4 Req 6B,200 crued wages payable 3,400 2, 900 Accounts receivable, net Notes receivable (trade) 2B, B00 Income taxes payable 5,000 36,150 Long-term note payable, secured by 71,400 mortgage on plant assets Prepaid expenses Plant assets, net 2, 950 Coamon stoch 152, 300 Retained earnings B1,000 63, 200 $ 247,400 Total assets $ 247.400 Total liabilities and equity These are short-term notes recelvable arising from customer (trade) sales. Required: Compute the following: (1) current ratio, (2) acld-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory. (6) debt-to-equity ratlo, (7) times Interest earned, (8) profit margln ratio. (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equity. (Do not round intermediate calculations.) Complete this question by entering your answers in the tabs below. Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the debt-to-equity ratio. Choose Numerator: Ratio otal liabilities Total equity ratio 2017 0 to 1 K Req 5 Req 7Notes receivable (trade) 5,000 36,150 Long-term note payable, secured by 71,400 mortgage on plant assets Prepaid expenses Plant assets, net 2, 950 Coamon stoch 152, 300 Retained earnings B1,000 63, 200 $ 247,400 Total assets $ 247.400 Total liabilities and equity These are short-term notes recelvable arising from customer (trade) sales. Required: Compute the following: (1) current ratio, (2) acld-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory. (6) debt-to-equity ratlo, (7) times Interest earned, (8) profit margln ratio. (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equity. (Do not round intermediate calculations.) Complete this question by entering your answers in the tabs below. Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the times interest earned. imes Interest Ear Choose Numerator: Choose - Times Interest Earned Income before taxnterest expense Interest expense Times interest earned 2017 0 times Req 8Total assets $ 247.400 Total liabilities and equity $ 241,400 These are short-term notes recelvable arising from customer (trade) sales. Required: Compute the following: (1) current ratio, (2) acld-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory. (6) debt-to-equity ratlo, (7) times Interest earned, (8) profit margln ratio. (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equity. (Do not round intermediate calculations.) Complete this question by entering your answers in the tabs below. Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req Req 9 Req 10Req 11 Compute the profit margin ratio. Choose Numerator: -Profit margin ratio ncome t sales Profit margin ratio 2017 < Req 7 Req 9Prepaid expenses Plant assets, net Total assets 2, 950 Coamon stoch 152, 300 Retained earnings B1,000 63, 200 $ 247,400 $ 247.400 Total liabilities and equity These are short-term notes recelvable arising from customer (trade) sales. Required: Compute the folloing: (1) current ratio, (2) acid-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory (6) debt-to-equity ratlo, (7) times Interest earned, (8) profit margin ratio, (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equity. (Do not round intermediate calculations.) Complete this question by entering your answers in the tabs below. Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the total asset turnover. l Asset Choose Numerator: I - Total Asset Turnover t sales Average total assets Total asset turnover 2017 0 times K Req 8 Req 10Total assets $ 247.400 Total liabilities and equity 5 247,400 These are short-term notes recelvable arising from customer (trade) sales. Required: Compute the following: (1) current ratio, (2) acld-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory. (6) deb on common stockholders equity. (Do not round Intermediate calculations.) ratio, (7) times Interest earned, (8) profit margin ratio, (9) total asset turnover, (10) return on total assets, and (11) return Complete this question by entering your answers in the tabs below. Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the return on total assets. (10) Return on Total Choose Numerator: I ncome Average total assets Return on total assets 2017 K Req 9 Req 11Plant assets, net 152,300 Retained earnings $ 247.400 Total liabilities and equity 63, 200 $ 247,400 Total assets These are short-term notes recelvable arising from customer (trade) sales. Required: Compute the following: (1) current ratio, (2) acld-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory. (6) debt-to-equity ratlo, (7) times Interest earned, (8) profit margln ratio. (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equity. (Do not round intermediate calculations.) Complete this question by entering your answers in the tabs below. Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the return on common stockholders equity. Return on Common Stockholders E Return On Common Stockholders Choose Numerator: Choose income- Preferred dividendsAverage common stockholders equityReturn on common stockholders equity 2017 K Req 10 Req 11
Answer #1

1 . Current ratio = current assets / current liabilities

= 92150/25800 = 3.57

Note

1 . Calculation of total current assets

Particulars Amount
Cash $14000
Short term investment $8200
Accounts receivable $28800
Notes receivable $5000
Merchandise inventory $36150
Total current assets $92150

* Please note that prepaid expense is not being included in the list of current asset because in the balance sheet it was separately mentioned from other current assets alongwith long term assets it was mentioned . If answer do not comes correct by using this current assets amount of $92150, then try it by adding the amount of prepaid expense amounting to $2950 and thus taking current assets to be $92150+$2950 = $95100.

2. Calculation of total currentl liabilities

Particulars amount
Accounts payable $19500
Accrued wages payable $3400
Income tax payable $2900
Total current liabilities $25800

2 . Quick ratio = quick assets / current liabilities

92150-36150/25800 = 2.17

3. Days sales uncollected = (accounts receivable + notes receivable /sales ) x 365

= (28800+5000/448600)x365 = 27.50

4 . Inventory turnover = COGS/average inventory

= 298150/{(36150+53900)/2} = 6.62

5. Days sales in inventory =( inventory / COGS)x364

= (36150/298150)x365 = 44.26 days

6 . Debt to equity ratio = total liabilities / total equity

= 97200/150200 = 0.65

7 . Times interest earned = (income before tax + interest)/ interest

= 47050+4400/4400= 11.69

8 . Profit margin ratio = net income / net sales x 100

= 28096/448600 x 100 = 6.26%

9 . Total assets turnover = net sales / average total assets

= 448600/{(247400+239400)/2}= 1.84

10 . Return on total assets = net income / average total assets = 28096/{(247400+239400)/2} = 11.54%

11 . Return on common stockholder equity = net income - preferred stock dividend / average common stockholder equity

= (28096-0)/{(150200+122104)/2} =20.64%